Exercise 22-18 (Algorithmic) (LO. 6) Noelle, the owner of all of the shares of ClockCo, an S corporation, transfers her stock to Grayson on April 1. ClockCo reports a $117,000 NOL for the entire tax year, but only $17,550 of the loss occurs during January-March Assume 365 days in a year. Round per day computations to five decimal places. Round final answers to the nearest dollar. Without a short-year election, $ of the loss is allocated to Noelle, and is allocated to Grayson If the corporation makes the short-year election, Grayson is allocated of the loss.
Answer #1
Without Short Year Election
Noelle 28849.31507 $
Grayson 88150.68493 $
With Short Year Election
Grayson 99450 $
Answer #2

Noelle's loss allocation: $17452 = 91:365 = x 70000

Grayson's loss allocation: $52548 = (365-91) = x 70000

If the corporation makes the short-year election, how much of the loss is allocated to Grayson?
$60000 = 70000-10000

answered by: anonymous

> Exercise 12-15 (LO. 6)

Noelle, the owner of all of the shares of ClockCo, an S corporation, transfers her stock to Grayson on April 1. ClockCo reports a $70,000 NOL for the entire tax year, but only $10,000 of the loss occurs during January-March.

Assume 365 days in a year. Round per day computations to five decimal places. Round final answers to the nearest dollar.

Without a short-year election, how much of the loss is allocated to Noelle and how much is allocated to Grayson?

Olik Le Sun, Mar 27, 2022 6:34 PM